# SAEmnesia: Erasing Concepts in Diffusion Models with Sparse Autoencoders

### v1

@@ 3 Methodology
Feature Centralization. As defined in Equation 5, feature centralization occurs when only one feature achieves a high score for a specific concept. However, computing the score function requires the entire dataset, making it impractical during training. Instead, we use $\mathcal{L}_{supSAE}$ as a proxy to achieve feature centralization. $\mathcal{L}_{CA}$ directly enforces that assigned latents activate strongly for their target concepts, while $\mathcal{L}_{L1}$ maintains sparsity across all latents. This combination concentrates a concept’s information in the designated latent.
Inference and Concept Unlearning. To unlearn a concept $c$, we first identify the single top SAE feature associated with $c$ and compute its average activation on a validation dataset $D$:
$F_{c}:=\{i|i\in\text{Top-}1(\{\text{score}(i,t,c,D)\})\}$ (11)
$\mu(i):=\mu(i,t,D),\forall i\in F_{c}$ (12)
@@
Note that this step is similar to the one proposed by Cywiński and Deja [9]; however, SAEmnesia only needs a single latent to erase each concept. The activation of the selected feature $F_{c}$ is multiplied by a negative value $\gamma_{c}<0$ normalized by the average activation on concept samples $\mu(i,t,D_{c})$. This removes the influence of the targeted concept on the activation vector $\mathbf{z}$. Each $i$-th latent feature activation is modified as follows:
$z_{i}=\begin{cases}\gamma_{c}\mu(i,t,D_{c})z_{i},&\text{if }i\in F_{c}\land z_{i}>\mu(i,t,D)\\ z_{i},&\text{otherwise}\end{cases}$ (13)
The condition $z_{i}>\mu(i,t,D)$ prevents random feature ablation when scores are low.
4 Experiments
We conduct comprehensive experiments to evaluate SAEmnesia across multiple dimensions: unlearning effectiveness, generation quality, concept separation, performance robustness, incremental unlearning capabilities, and adversarial resilience. Note that our focus here is on erasing objects, since it’s way more challenging that erasing styles [9]. This comprehensive evaluation framework allows us to demonstrate not only the effectiveness of our approach in removing unwanted concepts but also its practical applicability, robustness, and scalability in real-world deployment scenarios.
@@ 4 Experiments
Dataset. We extract internal activations from cross-attention blocks of SD1.5 [33] from UnlearnCanvas Benchmark [43]. Our approach differs from unsupervised methods in the data collection methodology through systematic concept labeling during data generation. Rather than requiring post-hoc concept discovery, we construct labeled training data by systematically generating activations from structured prompts with known concept compositions. For each object class $c$ (e.g., Bears, Cats) and style $s$ (e.g., Impressionism, Cubism), we generate 80 prompts of the form “a [object] in [style] style” and collect feature maps from selected U-Net cross-attention blocks across all 50 denoising timesteps during text-conditioned generation. Each activation is directly labeled with its corresponding object and style based on the prompt structure, creating explicit concept-activation pairs. The employed objects and styles are taken from the UnlearnCanvas benchmark, which consists of 20 different objects and 50 different styles.
We focus our analysis on block up.1.1 for object-related features, as this block has been empirically demonstrated to specialize in generating specific visual aspects [2]. This controlled labeling strategy provides clean supervision signals that enable direct concept-neuron mapping during SAE training, contrasting with unsupervised approaches that must discover concept representations through post-hoc analysis. For evaluation, we employed the same setting of SAeUron [9].
@@
SAE Model. Our best-performing method starts with a pre-trained unsupervised SAE and fine-tunes it using CA loss (8) combined with the orthogonality constraint (9) to enforce concept separation. The weights chosen for the loss function components and additional model setups can be found in the Appendix 7.107.57.6 and 7.7).
@@
Evaluation Metrics. We employ a comprehensive evaluation framework that assesses multiple aspects of unlearning performance and model behavior. Our primary evaluation uses the UnlearnCanvas benchmark [43] with Vision Transformer-based classifiers to measure three key metrics: Unlearning Accuracy (UA), which quantifies the proportion of samples from target concept prompts that are misclassified (i.e., successful unlearning); In-domain Retain Accuracy (IRA), measuring classification accuracy on retained concepts within the same domain; and Cross-domain Retain Accuracy (CRA), assessing accuracy on concepts from different domains such as object accuracy during style unlearning. To evaluate generation quality, we compute Fréchet Inception Distance (FID) scores across different unlearning configurations and multiplier values, quantifying both the quality and diversity of generated images. Complete results are reported for each tested model in Table 7 and Table 8 in the Appendix. We also conduct specialized concept separation analysis by examining latent overlap between different concepts, with particular focus on the challenging Dogs vs. Cats classification case, a known limitation of current SAE-based unlearning approaches [9] that leads to concept interference, Results are presented in Appendix 7.8.
Our robustness evaluation also includes systematic analysis of model performance across different uniform multiplier values, providing insights into stability and tunability across various unlearning intensities. We demonstrate incremental unlearning capability through sequential object removal experiments, showing that our method can handle dynamic unlearning scenarios without catastrophic interference.
For adversarial robustness, we follow Zhang et al. [42] and evaluate against UnlearnDiffAtk attacks, optimizing 5-token adversarial prefixes for 40 iterations with learning rate 0.01 to provide fair comparison with baseline SAE methods [9]. Additionally, we provide UA-IRA confusion matrices for intuitive visualization of trade-offs between unlearning effectiveness and retention accuracy across different experimental conditions.
5 Results
We evaluate our supervised SAE training approach across multiple model variants, exploring both training from scratch and fine-tuning existing checkpoints. Among the several tested configurations, we present the detailed analysis of SAEmnesia as our top performer. Additional results for tested variants are available in the Appendix 7.7.
Concept-Latent Association Distribution. Our core contribution is creating interpretable one-to-one concept-neuron mappings. Figure 3 is an example of the effectiveness of our supervised training in promoting feature centralization by comparing feature importance score distributions before and after training for the Flowers concept. Additional examples of score distributions are reported in the Appendix 7.3. The higher score values are also tightly linked to the lower absolute values of the optimal unlearning multipliers (a deeper analysis can be found in Appendix 7.2).
[…]
The original SAeUron model exhibits relatively uniform, low-magnitude scores across the entire latent space (max score: 0.015), indicating distributed concept representation. Our supervised SAEmnesia model produces a clear dominant peak at neuron 11979 with a maximum score of 0.04, 2.67 times higher than the baseline. This concentrated activation pattern confirms that supervised training successfully enforces strong one-to-one concept-latent relationships, with the assigned neuron becoming highly specialized for the target concept while other neurons remain largely inactive.
This transformation from distributed to concentrated representations is the foundation that enables all subsequent improvements in computational efficiency and unlearning performance.
@@ 5 Results
UnlearnCanvas Benchmark Performance. Table 1 presents our performance compared to the SAeUron baseline. An inconsistency in the computation of the metrics was discovered in the public implementation of SAeUron. We therefore recomputed the metrics ourselves using the corrected procedure. SAEmnesia achieves 91.51% average score with hyperparameter search, improving over the baseline’s 82.29%. Standard deviations are reported in Appendix 7.1.
@@
ESD [14] 92.15 55.78 44.23 64.05
FMN [41] 45.64 90.63 73.46 69.91
UCE [15] 94.31 39.35 34.67 56.11
CA [24] 46.67 90.11 81.97 72.92
SEOT [26] 23.25 95.57 82.71 67.18
SPM [28] 71.25 90.79 81.65 81.23
EDiff [40] 86.67 94.03 48.48 76.39
SHS [39] 80.73 81.15 67.99 76.62
SAeUron [9] 87.16 85.57 74.14 82.29
@@
The performance gains stem from the stronger concept-latent associations that enable more targeted interventions. Therefore, SAEmnesia is particularly effective for maintaining concept separation.
@@
Effect of Uniform Multipliers. Figure 4 compares the effect of applying uniform multipliers (so the same multiplier for all objects) to the latent steering for SAeUron and SAEmnesia. Across all evaluation metrics, SAEmnesia consistently achieves higher and more stable scores over the full range of multipliers. This demonstrates that SAEmnesia is less sensitive to the exact steering strength across different objects and provides more robust control of concept removal. Additional quantitative plots are included in the Appendix 7.9.
@@
Figure 4: Effect of uniform multipliers on unlearning performance for SAeUron and SAEmnesia. SAEmnesia maintains higher and more stable scores across all multipliers compared to SAeUron, indicating greater robustness to the steering strength.
@@
Qualitatively Results. Figure 5 shows qualitative examples of SAEmnesia applied to a representative subset of object–style combinations. Thanks to the one–to–one mapping between concepts and latent units, steering the sparse autoencoder selectively removes the targeted concept: the removed concepts vanish in the diagonal images, while remaining visible in the off–diagonal ones. For highly entangled concepts such as Human, some residual shapes or textures may still appear after removal, reflecting the greater difficulty of isolating and suppressing such features. Additional qualitative examples covering a wider range of objects and styles are provided in Appendix 7.4.
@@
Sequential Unlearning Scalability. To demonstrate how interpretable representations enable scalability, we test sequential unlearning across 9 objects: Bears $\rightarrow$ Bears+Cats $\rightarrow\rightarrow$ 9 total objects. This challenging scenario tests whether one-to-one mappings maintain effectiveness as complexity increases.
Table 2 shows that SAEmnesia significantly outperforms the baseline. SAEmnesia achieving 92.4% accuracy for 9-object removal compared to the baseline’s 64%. This scalability directly results from having interpretable, specialized neurons rather than distributed representations that interfere with each other.
Table 2: Evaluation of SAEmnesia variants against SAeUron baseline on sequential object unlearning tasks. The best result for each task is highlighted in bold. SAEmnesia 92.4% accuracy on 9-object removal compared to SAeUron’s 64%, representing a 28.4% improvement.
Task SAeUron [9] SAEmnesia
0 (Bears) 0.98 0.99
1 (Bears+Cats) 0.56 1.0
2 (3 objects) 0.45 1.0
3 (4 objects) 0.48 0.9975
4 (5 objects) 0.48 1.0
5 (6 objects) 0.58 0.895
6 (7 objects) 0.61 0.91
7 (8 objects) 0.63 0.916
8 (9 objects) 0.64 0.924
However, this comes with retention accuracy trade-offs (Table 3). Importantly, all models maintain their operational ranges throughout the sequential unlearning process. SAEmnesia demonstrates improved retention performance, achieving 60.9% final retention compared to the baseline’s 48.4%.
Table 3: Sequential Unlearning - Retention Accuracy Comparison
Model RA Range Final RA
SAeUron [9] 40-48% 48.4%
SAEmnesia 46-61% 60.9%
@@
Adversarial Robustness. The concentrated concept representations also improve adversarial robustness, as shown in Figure 6. Under UnlearnDiffAtk attacks, SAEmnesia demonstrates improved resilience, dropping from 73.30% to 28.30% unlearning accuracy. This contrasts sharply with the state-of-the-art significant drop from 26.19% to 2.22%. The supervised concept-neuron mapping approach in SAEmnesia shows substantially improved robustness against adversarial perturbations, with the stronger concept-neuron bonds making it significantly harder for attacks to disrupt the specialized representations.

### v2

@@ 3 Methodology
Feature centralization. As defined in Eq. 5, feature centralization occurs when only one feature achieves a high score for a specific concept. However, computing the score function requires the entire dataset, making it impractical during training. Instead, we use $\mathcal{L}_{\text{supSAE}}$ as a proxy to achieve feature centralization. $\mathcal{L}_{\text{CA}}$ directly enforces that assigned latents activate strongly for their target concepts, while $\mathcal{L}_{\text{L1}}$ maintains sparsity across all latents. This combination concentrates a concept’s information in the designated latent.
@@
Inference and concept unlearning. To unlearn a concept $c$, SAEmnesia only needs the single latent $i_{c}$ to erase the concept. The activation of the selected feature is multiplied by a negative value $\gamma_{c}<0$ normalized by the average activation $\mu(i_{c},t,D_{c})$ on concept samples of a validation dataset $D$. This removes the influence of the targeted concept on the activation vector $\mathbf{z}$. Each $i_{c}$-th latent feature activation is modified as follows:
$z_{i_{c}}=\begin{cases}\gamma_{c}\mu(i_{c},t,D_{c})z_{i_{c}},&\text{if }z_{i_{c}}>\mu(i_{c},t,D)\\ z_{i_{c}},&\text{otherwise}\end{cases}$ (10)
The condition $z_{i_{c}}>\mu(i_{c},t,D)$ prevents random feature ablation when scores are low. During inference, we can use the original pretrained model for the first $t$ steps, setting the multipliers to 1 and retaining the pretrained model priors. We can then turn on SAEmnesia for the remaining steps.
4 Experiments and Results
We conduct comprehensive experiments to evaluate SAEmnesia across multiple dimensions: unlearning effectiveness, generation quality, concept separation, performance robustness, incremental unlearning capabilities, adversarial resilience and nudity removal. Note that our focus in this work is on object erasure, as it presents a considerably greater challenge than style erasure [cywinski2025saeuron].
4.1 Experimental setup
We report here the evaluation setup for the experiments of our proposed method, including datasets, architectures and evaluation metrics.
@@ 4.1 Experimental setup
Dataset. We extract activations from SD v1.5 [rombach2022high] within the UnlearnCanvas Benchmark [zhang2024unlearncanvas]. SD v1.x family remains the standard evaluation setting for concept-erasure research: the benchmark itself is built on SD v1.5, and most contemporary methods report results on SD v1.x [cywinski2025saeuron, Thakral_2025_CVPR, Wang_2025_CVPR, ACE_2025_CVPR, Lee_2025_CVPR, Srivatsan_2025_CVPR]. This makes SD v1.x the appropriate and widely-accepted testbed for fair comparison and reproducibility. We construct labeled training data by generating activations from structured prompts with known concept compositions. For each object class $c$ (e.g., “Bears”, “Cats”) and style $s$ (e.g., “Impressionism”, “Cubism”), we generate 80 prompts of the form An image of {object} in {style} and collect feature maps from selected U-Net cross-attention blocks across all 50 denoising timesteps during text-conditioned generation. Each activation is directly labeled with its corresponding object and style based on the prompt structure, creating explicit concept-activation pairs. The employed objects and styles are taken from the UnlearnCanvas benchmark, which consists of 20 different objects and 50 different styles. We focus our analysis on block up.1.1 for object-related features, as this block has been empirically demonstrated to specialize in generating specific visual aspects [basu2023localizing]. This controlled labeling strategy provides clean supervision signals that enable direct concept-neuron mapping during SAE training, contrasting with unsupervised approaches that must discover concept representations through post-hoc analysis. For evaluation, we employed the same setting of SAeUron [cywinski2025saeuron].
@@
SAE model. Our best-performing method starts with a pre-trained unsupervised SAE and fine-tunes it using SAEmnesia loss in Eq. 6. Unless otherwise stated, SAEmnesia is applied across all denoising steps for a fair comparison with [cywinski2025saeuron]. For the decorrelation constraint in Eq. 9, we choose $\mathcal{C}_{\text{obj}}$ and $\mathcal{C}_{\text{sty}}$, corresponding to object and style concepts, to enforce concept separation. Loss function hyperparameters and additional model setups can be found in the Appendix 6.56.66.7 and 6.10.
@@
Evaluation metrics. Our primary evaluation uses the UnlearnCanvas benchmark [zhang2024unlearncanvas] with Vision Transformer-based classifiers to measure three key metrics: Unlearning Accuracy (UA), which quantifies the proportion of samples from target concept prompts that are misclassified (i.e., successful unlearning); In-domain Retain Accuracy (IRA), measuring classification accuracy on retained concepts within the same domain; and Cross-domain Retain Accuracy (CRA), assessing accuracy on concepts from different domains such as object accuracy during style unlearning. To evaluate generation quality, we compute Fréchet Inception Distance (FID) scores across different unlearning configurations and multiplier values, quantifying both the quality and diversity of generated images.
4.2 Quantitative results
@@
UnlearnCanvas benchmark performance. Tab. 1 presents our performance compared to the SAeUron baseline. An inconsistency in the computation of the metrics was discovered in the public implementation of SAeUron. We therefore recomputed the metrics ourselves using the corrected procedure. SAEmnesia achieves 91.51% average score with hyperparameter search, improving over the baseline’s 82.29%. Standard deviations are reported in Appendix 6.1. The performance gains stem from the stronger concept-latent associations that enable more targeted interventions. Therefore, SAEmnesia is particularly effective for maintaining concept separation. The performance gains stem from the stronger concept-latent associations that enable more targeted interventions.
@@
ESD [gandikota2023erasing] 92.15 55.78 44.23 64.05
FMN [zhang2024forget] 45.64 90.63 73.46 69.91
UCE [gandikota2024unified] 94.31 39.35 34.67 56.11
CA [kumari2023ablating] 46.67 90.11 81.97 72.92
SEOT [li2024get] 23.25 95.57 82.71 67.18
SPM [lyu2024one] 71.25 90.79 81.65 81.23
EDiff [wu2024erasediff] 86.67 94.03 48.48 76.39
SHS [wu2024scissorhands] 80.73 81.15 67.99 76.62
SAeUron [cywinski2025saeuron] 87.16 85.57 74.14 82.29
@@
We also conduct specialized concept separation analysis by examining latent overlap between different concepts, with particular focus on the challenging “Dogs vs. Cats” classification case, a known limitation of current SAE-based unlearning approaches [cywinski2025saeuron] that leads to concept interference. Results are presented in Appendix 6.8.
@@
Effect of uniform multipliers. Fig. 3 compares the effect of applying uniform multipliers (so the same multiplier for all objects) to the latent steering for SAeUron and SAEmnesia. Across all evaluation metrics, SAEmnesia consistently achieves higher and more stable scores over the full range of multipliers. This demonstrates that SAEmnesia is less sensitive to the exact steering strength across different objects and provides more robust control of concept removal. Additional quantitative plots are included in the Appendix 6.9.
@@
Figure 3: Effect of uniform multipliers on unlearning performance for SAeUron and SAEmnesia. SAEmnesia maintains higher and more stable performance across all multipliers compared to SAeUron, indicating greater robustness to the steering strength.
Concept-latent association distribution. Our core contribution is creating interpretable one-to-one concept-neuron mappings. Fig. 4 is an example of the effectiveness of our supervised training in promoting feature centralization by comparing feature importance score distributions before and after training for the “Flowers” concept. The original SAeUron model exhibits relatively uniform, low-magnitude scores across the entire latent space (max score: 0.0166), indicating distributed concept representation. Our supervised SAEmnesia model produces a clear dominant peak at neuron 11979 with a maximum score of 0.0404, 2.43 times higher than the baseline. This concentrated activation pattern confirms that supervised training successfully enforces strong one-to-one concept-latent relationships, with the assigned neuron becoming highly specialized for the target concept while other neurons remain largely inactive. This transformation from distributed to concentrated representations is the foundation that enables all subsequent improvements in computational efficiency and unlearning performance. The higher score values are also tightly linked to the lower absolute values of the optimal unlearning multipliers (a deeper analysis can be found in Appendix 6.2). Additional examples of score distributions are reported in the Appendix 6.3.
[…]
Feature centralization validation via K-NN classification. To quantitatively validate that SAEmnesia achieves feature centralization, where each concept’s information is concentrated in a single latent, we conducted k-nearest neighbors (k-NN) classification experiments on the latent representations across all 20 object concepts in the UnlearnCanvas benchmark. For each object, we compared classification accuracy across the denoising process using four strategies: (1) the top-scoring latent identified by SAEmnesia (Eq. 4), (2) all latent features, (3) randomly selected features, and (4) random guess baseline. As shown in Fig. 5, the score-based selection achieves nearly identical classification accuracy to using all available latents throughout most of the denoising timesteps.
Figure 5: K-NN classification across denoising timesteps, averaged over 20 object concepts. Using only the top-scoring latent identified by SAEmnesia, performances are similiar to using all features, demonstrating that supervised training successfully concentrates concept-relevant information into single interpretable latents across diverse object categories.
4.3 Qualitative results
@@
Concepts removal. Figs. 1 and 6 show qualitative examples of SAEmnesia applied to a representative subset of object–style combinations. In this setting, we use concept multipliers of $\gamma_{c}=-1$ and restrict the application of SAEmnesia to the final 25 denoising steps to reduce artifacts. Thanks to the one–to–one mapping between concepts and latent units, steering the sparse autoencoder selectively removes the targeted concept: the removed concepts vanish in the diagonal images, while the corresponding style is preserved. When unlearning unrelated content, the objects in the prompts remain present in the non-diagonal ones. Additional qualitative examples covering a wider range of objects and styles are provided in Appendix 6.4.
@@
SAEmnesia effect on discovered concepts. We examine the activation of the concept selected feature on corresponding image patches across two objects, “Architectures” and “Rabbits”. Fig. 7 shows the difference between unsupervised SAE training and our SAEmnesia approach. In the unsupervised setting (first rows), we can see that unsupervised SAEs decompose concepts into multiple distributed features of comparable importance. In contrast, SAEmnesia (second rows) produces a markedly different activation structure. This concentrated representation, where object-level semantics are encoded in a single, highly interpretable latent, facilitates more precise concept manipulation and enables effective machine unlearning through targeted interventions.
Figure 7: SAEmnesia shifts attention toward patches most responsible for the target concept. Visualization of the most important patches for the objects “Architectures” (top) and “Rabbits” (bottom) across timesteps.
4.4 Additional experiments
Sequential unlearning scalability. To demonstrate how interpretable representations enable scalability, we test sequential unlearning (where unlearning requests arrive sequentially) across 9 objects: “Bears”, “Cats”, “Flowers”, “Frogs”, “Jellyfish”, “Sea”, “Statues”, “Sandwiches” and “Waterfalls”. We selected these objects because the UnlearnCanvas benchmark focuses only on the sequential unlearning of styles. In this setting, we apply the unlearning multipliers cumulatively: first $\gamma_{\text{Bears}}$ alone, then $\gamma_{\text{Bears}}$ and $\gamma_{\text{Cats}}$, and so on. Fig. 8 shows that SAEmnesia significantly outperforms the baseline. SAEmnesia achieves 92.4% accuracy for 9-object removal compared to the baseline’s 64%. This scalability results from having interpretable, specialized neurons rather than distributed representations that interfere with each other. SAEmnesia also achieves higher retention accuracy (RA, average of IRA and CRA), reflecting the model’s ability to preserve all non-removed concepts. When removing all objects, SAEmnesia attains RA of 60.9%, while the baseline 48.4%.
Figure 8: Evaluation of SAEmnesia against SAeUron baseline on sequential object unlearning tasks. SAEmnesia achieves higher UA and RA.
@@ 4.4 Additional experiments
Adversarial robustness. We follow zhang2024generate and evaluate against UnlearnDiffAtk attacks, optimizing 5-token adversarial prefixes for 40 iterations with learning rate 0.01 to provide fair comparison with baseline SAE methods [cywinski2025saeuron]. The concentrated concept representations also improve adversarial robustness, as shown in Fig. 9. SAEmnesia demonstrates improved resilience, dropping from 73.30% to 28.30% unlearning accuracy. This contrasts sharply with the state-of-the-art significant drop from 26.19% to 2.22%. The supervised concept-neuron mapping approach in SAEmnesia shows substantially improved robustness against adversarial perturbations, with the stronger concept-neuron bonds making it significantly harder for attacks to disrupt the specialized representations.